Fermi National Laboratory
 

Operation Summary for April 29, 2011 - May 2, 2011

 

The Integrated Luminosity for the period from 4/25/11 to 5/2/11 was 71.36 pb-1 (inverse picobarns). NuMI took 4.06E18 of beam during this same time period.

Friday April 29

The day shift began with Operations monitoring Tevatron (TEV) store 8700 (76.11E30) and the Recycler stash (512.1E10), with the Antiproton Source (Pbar) stacking (19.8E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations terminated store 8700 at 9:52 AM, put the TEV into HEP shot setup, and established store 8701 at 11:04 AM with an average initial luminosity (AIL) of 416.61E30. Store 8701 set a new proton-antiproton initial luminosity record. MI Personnel conducted studies in the afternoon.

During the evening shift, Operations disabled beam to T-978 at 6:04 PM. Operations reported some short downtimes due to Booster RF system trips.

Saturday April 30

The midnight shift began with Operations monitoring the store 8701 (101.33E30) and the stash (324.5E10), with Pbar stacking (24.2E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to T-978 at 4:50 AM. Operations terminated store 8701 at 5:20 AM, put the TEV into shot setup, and established store 8702 at 6:48 AM with an AIL of 405.1E30.

During the day shift, Operations reported that module problems with a Pbar Debuncher kicker (D:IKIK) interrupted stacking for 3 hours.

During the evening shift, Operations disabled beam to T-978 at 6:05 PM.

Sunday May 1

The midnight shift began with Operations monitoring store 8702, the stash, and the Pbar stack, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8702 at 4:28 AM, put the TEV into shot set up, and established store 8703 at 5:57 AM with an AIL of 402.91E30. Operations reported that a beam instability occurred at the beginning of store 8703 that caused a 50E30 loss of luminosity.

During the day shift, Operations reported that Booster RF system trips cause minimal downtime during the shift.

During the evening shift, Operations disabled beam to T-978 at 6:05 PM. Operations terminated store 8703 at 9:35 PM and prepared D0 and CDF for accesses.

Monday May 2

The midnight shift began with D0 and CDF in access, with Operations the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, D0 and CDF were both out of access by 12:04 AM. Operations put the TEV into shot set up and established store 8707 at 3:33 AM with an AIL of 408.95E30.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow MI personnel to conduct a coalescing study and a g-2 study. Depending on circumstances, he might also allow a TEV access at the end of store 8707 for various maintenance jobs.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/02/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory